OpenStudy (anonymous):

will medal Express your answer in lowest terms 1/6 x 5/7

OpenStudy (kitten_is_back):

1/6 = 7/42 then you have to reduce it to 1/6 and 5/7 = 30/42 and when you reduce it 5/7 it think